OpenStudy (anonymous):

Find the length of the hypotenuse of an isosceles right triangle whose legs are 1 unit in length.

OpenStudy (anonymous):

isosceles right triangle means that both of the legs are 1 unit in length so use pythagoran theorem. \[c = \sqrt{a^2 + b^2}\]

OpenStudy (anonymous):

c = \[\sqrt{1^{2 }+ 1^{2}}\]

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

so c = 1 + 1 c = 2

OpenStudy (anonymous):

\[c = \sqrt{a^2 + b^2}\]

OpenStudy (anonymous):

so c = \[\sqrt{1 + 1} = \sqrt{2}\]???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

you're welcome